Let A=left−1,1,−i,iright,i=sqrt−1. Let the binary operation be multiplication, what is the inverse element of i?
a.\\(-i\\)
b.\\(i\\)
c.1
d.\\(-1\\)
1
Expert's answer
2020-05-20T20:00:22-0400

Given "A= \{1,-1,i,-i\}" and binary operation ia multiplication.

Identity of given set is 1 under binary operation multiplication because "a \times 1 = 1\times a = a \ \forall \ a\in A".

Inverse of "a\in A" is "b" if "b \in A" and "a\times b = b \times a = 1".


So, Inverse of "i" is "-i" because "i \times (-i) = (-i) \times i = -i^2 = 1" and also "-i \in A".

So, Answer is option a) "-i".